To determine the number
of ounces of fluid that a person should drink in a
day, divide his or her weight in pounds by 2 and then
add 0.4 ounce for every minute of exercise.
(a) Write a function that gives the fluid requirements
for a person weighing 150 pounds and exercising
x minutes a day.
(b) If a 150-pound runner needs 89 ounces of fluid
each day, determine the runner’s daily minutes of
exercise.

Answer by John10(297)

a) Let x be the number of daily minutes and f(x) is the function (number of ounces of fluid):
f(x) = w/2 + 0.4x where w is the weight.
If w = 150 lbs then
f(x) = 150/2 + 0.4x = 75 + 0.4x
b) If the f(x) = 89 oz
0.4x + 75 = 89
0.4x = 89 - 75
0.4x = 14
x = 14/0.4
x = 35 oz of fluid
John10:)